OpenStudy (anonymous):

integrate (x)(cos^3 x) dx

OpenStudy (sirm3d):

do you know the reduction formula for \[\int \cos^n x \,\mathrm dx\]

OpenStudy (anonymous):

reduction?? nope...

OpenStudy (anonymous):

first simplify cos^3 x in terms of cosine of 3x and x and then integrate by parts

OpenStudy (anonymous):

i don't think here u should use reduction formula

OpenStudy (anonymous):

its \[\cos ^3 x\] .. where did u get 3x and x @matricked

OpenStudy (anonymous):

use cos3x =4 cos^3 (x) - 3cos x

OpenStudy (sirm3d):

why not? it solves the problem too, but the solution may not be attractive to @bii17

OpenStudy (anonymous):

i mean use cos^3 (x) =1/4 *(cos3x +3cosx)

OpenStudy (anonymous):

oh okay.. then what will happen next ? @matricked

OpenStudy (anonymous):

and then use by parts

OpenStudy (anonymous):

\[1/4 [x(\cos 3x + 3\cos x)]dx\] is this correct? @matricked

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so there will be u and du den dv and v... i used Dv as cos 3x dx the V will be 1/3 sin 3x ???

OpenStudy (anonymous):

for both integration of 1/4 *xcos3x take x as the first function

OpenStudy (anonymous):

i get it.. thanks a lot :) @matricked